# The Generalization Ridge: Information Flow in Natural Language Generation

### v2

@@ 7 Conclusion
We introduce InfoRidge, an information-theoretic framework designed to trace and quantify how information evolves across layers in transformer-based language models for natural language generation. By estimating both predictive information and incremental information gain, we systematically characterize the layerwise dynamics of information flow, offering a principled view of how signals are refined, amplified, or diminished as they propagate through the network. Our findings reveal a consistent generalization ridge emerging in intermediate layers, where mutual information between the hidden representation and the target label reaches its peak before gradually declining. This phenomenon reflects a fundamental trade-off between generalization and memorization as information flows deeper into the model. Residual scaling experiments further corroborate this interpretation, demonstrating the functional specialization of different layers. Taken together, these findings position InfoRidge as a comprehensive framework for diagnosing how language models internally manage information during natural language generation, while also revealing the structural mechanisms that govern the balance between generalization and memorization in transformer architectures.
Ethics Statement.
This work adheres to the Code of Ethics. Our experiments use only open-source models and publicly available datasets under their respective open licenses, with no involvement of human subjects or sensitive data. We identify no foreseeable ethical risks.
Reproducibility Statement.
We ensure reproducibility by providing experimental and implementation details in Section 4 and Appendices BC. Full results with statistical significance are in Appendix E, and anonymous source code is included as supplementary material.
@@ A.2 Mutual Information Estimation
A.3 Min Wasserstein Generalization Bound (He et al., 2024)
We restate the generalization bound proposed by He et al. (He et al., 2024), which characterizes generalization in terms of the Wasserstein distance between internal representations.
Suppose that the loss function $\tilde{\ell}:\mathcal{Y}\times\mathcal{Y}\to\mathbb{R}_{\geq 0}$ is $\rho_{0}$-Lipschitz, and the activation function $\phi_{\ell}:\mathbb{R}\to\mathbb{R}$ is $\rho_{\ell}$-Lipschitz for each $\ell=1,\ldots,L$. Then:
$\mathrm{gen}(P_{W|\mathcal{D}_{n}}\!,\!P_{X,Y})\!\leq\!\min_{\ell}\frac{\rho_{0}}{n}\sum_{i=1}^{n}\!\mathbb{E}_{W}\!\left[\!\left(1\!\vee\!\prod_{j=\ell+1}^{L}\!\rho_{j}\|W_{j}\|_{\mathrm{op}}\!\right)\!\cdot\!\mathcal{W}_{1}\!\left(\!P_{T_{\ell},i,Y_{i}|W}(\cdot|W),\!\ P_{T_{\ell},Y|W}(\cdot|W)\!\right)\!\right]$ (3)
This result shows that generalization can be tightly controlled by the Wasserstein distance (Villani et al., 2008) between representations at a specific layer—referred to the generalization funnel layer.
Connection to Our Work.
The Min Wasserstein Generalization Bound (He et al., 2024) provides a theoretical foundation for our study by characterizing generalization in terms of distributional alignment at an intermediate layer. It motivates our analysis of information flow by suggesting that information is peaked at a specific layer—the generalization funnel. Our InfoRidge builds on this insight by quantifying predictive information across layers, and reveals that a specific intermediate layer exhibits peak mutual information and correlates with better generalization performance.
@@ Appendix B Dataset Overview and Statistics
To evaluate information flow and generalization dynamics across model layers, we conduct experiments on three datasets with varying levels of complexity and structure: CLUTRR, ECQA, and a custom-designed Synthetic Arithmetic dataset. Table B.1 summarizes key dataset statistics.
B.1 Dataset Overview
@@ C.1 Prompt Construction
All tasks are cast into a next-token generation format. The model receives a prompt and generate the next token. Below are construction strategies and examples for each dataset:
@@ C.2 Finetuning Settings
C.3 Information Estimation
To estimate mutual information, we subsample between 50 and 200 test examples depending on model and task to achieve a stable result.
C.4 Residual Scaling with Learnable $\beta_{\ell}$ Parameters
@@ Appendix D Compute and Licensing Details
CLUTRR (Sinha et al., 2019) is released under a CC BY 4.0 license as part of EMNLP 2019. The ECQA (Aggarwal et al., 2021) dataset is also released under a CC BY 4.0 license by its authors as part of EMNLP 2021. The Synthetic Arithmetic dataset is custom-designed by the authors and does not rely on any external or licensed data sources.
@@ Appendix E Full Quantitative Results with Confidence Estimates
To provide a comprehensive view of model behavior, we report the full experimental results for all models and datasets. This includes predictive information trends ($I(Z;Y)$), incremental information gain ($I(\Delta Z;Y)$) and residual scaling effects ($\beta_{\ell}$), with statistical significance reporting.
E.1 Predictive Information
To estimate mutual information between hidden representations and target labels, we compute matrix-based mutual information following the methodology in Appendix A. We subsample between 50 and 100 test examples per experiment and report results with statistical significance based on multiple random seeds.
@@ E.1 Predictive Information
Figure 7: Predictive information $I(Z;Y)$ across different models and datasets exhibits an information peak, indicating a generalization ridge. In cases where the task is too simple relative to model capacity—such as the synthetic arithmetic task with LLaMA—this trend reflects an overfitting regime. Lighter line colors represent later training epochs. Each curve shows the mean across 5 random seeds (0, 1, 2, 3, 42), and the shaded region denotes a 2-sigma (~96%) confidence interval.
E.2 Incremental Information Gain
@@
Figure 8: Incremental information gain $I(\Delta Z;Y)$ across different models and datasets with ~96% CI error bars. Across all models, we observe that the largest incremental information gain consistently occurs in intermediate layers—further supporting the emergence of a generalization ridge.
E.3 Residual Scaling
We introduce a residual scaling mechanism with learnable scalar coefficient parameters, inspired by prior work on adaptive residual modulation (Liu et al., 2019a; Menghani et al., 2024). Similar ideas of modulating internal computation have also been explored in parameter-efficient fine-tuning (PEFT) (Houlsby et al., 2019), Representation-Efficient Fine-Tuning (REFT) (Wu et al., 2024) and interpretability-driven control (Huang et al., 2024; Deng et al., 2025; Meng et al., 2022; Wu et al., 2023). We present the complete set of residual scaling results, detailing the learned $\beta_{\ell}$ values across all transformer layers. These values reflect the relative contribution of each layer after optimizing the residual scaling coefficients while keeping all other model parameters frozen. $\beta_{\ell}$ are trained on in-distribution (ID) and out-of-distribution (OOD) data. We observe that in the ID setting, later layers tend to receive higher weights, consistent with memorization behavior. In contrast, OOD training consistently has lower final layers’ weight compared to the ID settings and shifts importance toward the intermediate layers. When the task is simple relative to the model’s capability, such as LLaMA on the synthetic task, the final-layer residual coefficients are similar for ID and OOD, consistent with Figure 7. The in-distribution (ID) setting is using the training split of “gen_train23_test2to10”. Out-of-distribution (OOD) is conducted on held-out CLUTRR configurations that do not overlap with the training distribution: for GPT models we evaluate on the test split of “gen_train23_test2to10” (lengths 2 through 10), while for Qwen and LLaMA we use the test split of “rob_train_sup_23_test_all_23”, which defines a more challenging OOD setting by jointly varying both the clause length and the status of noise facts introduced into the story, given the higher modeling capacity of these models. In all cases, tasks 1.2 and 1.3 are excluded from the OOD sets, as they overlap with the training configuration and are therefore treated as in-distribution rather than out-of-distribution.
@@ E.3 Residual Scaling
Figure 9: Residual scaling coefficients $\beta_{\ell}$ across all transformer layers. ID training emphasizes later layers, while OOD training has lower final layer weight. Each curve shows the mean across five random seeds (0, 1, 2, 3, 42), and the shaded region denotes 1-sigma error bar.
Appendix F Attention Dynamics Across Layers Case Study
To verify whether generalization ridge layer indeed correspond to semantically meaningful processing, we visualize the attention patterns as a more interpretable signal of where the model focuses. As shown in Figure 10, we visualize attention maps across layers.
Figure 10: Attention map across layers (GPT-2 Small, Synthetic). At the generalization ridge layers, attention becomes more targeted, focusing on predictive tokens.
In early layers (e.g., Layer 1), attention is diffuse and biased toward final tokens—reflecting reliance on position rather than true predict signals.
By generalization ridge layers (e.g., Layer 11), attention becomes more targeted, concentrating on predictive tokens. This shift marks a transition from positional attention to semantically meaningful focus, suggesting that intermediate layers are increasingly capable of isolating task-relevant information from irrelevant content.
In the final layers, attention regresses toward final tokens. This reversion is aligned with the observed decline in $I(Z_{\ell};Y)$. The resurgence of attention towards terminal tokens indicates a potential memorization scenario, where the model re-engages superficial positional strategies, possibly memorizing noise rather than further refining the generalized extraction of predictive signals.
This attention trajectory further supports the generalization ridge hypothesis, highlighting a trade-off between generalization and memorization in the model’s representational strategy.
Appendix G Incremental Information Gain Case Study
To further illustrate how intermediate layers contribute to generalization, we analyze the semantic content introduced by residual transitions at different depths. Specifically, we decode the residual transition $\delta z$ at each layer using the language modeling (LM) head, projecting the incremental representation back into token space.
This analysis allows us to inspect the linguistic shift introduced by each transformer block in isolation, and to assess whether the changes correspond to task-relevant predictions or superficial noises.
G.1 Methodology
For a fixed input, we compute residual transitions $\delta z^{(\ell)}$ at each layer and pass them through the model’s final linear projection (LM head) followed by a softmax. We record the top predicted tokens and their probabilities, and it reflects the directional change applied by layer $\ell$.
G.2 Observations
We decode residual transitions across layers and report the top shifted tokens by projecting $\Delta Z^{(\ell)}$ through the LM head. These shifts provide insight into how each layer modifies the model’s internal prediction trajectory.
Layer 9 → 10: The top shifted tokens include ĠNine, Ġ10, Ġ4, Ġ8, and 8, which are all numerically aligned with the target prediction space in the synthetic arithmetic task. This indicates that the model is beginning to refine task-relevant numerical features at this depth.
Layer 10 → 11: The shifted tokens become partially diluted, featuring punctuation and less informative symbols such as ,, Ġ, and ., alongside occasional task-relevant entries like Ġ12 and Ġ4. This indicates a transitional phase where the model continues to refine meaningful task-relevant features, yet begins to exhibit noise from frequent but semantically uninformative, such as punctuation, tokens.
Layer 11 → 12: At the final layer transition, the top 5 shifted tokens become largely uninterpretable, including Ġcanaver, soDeliveryDate, enegger, 76561, and ILA. Conversely, the most negatively shifted tokens—Ġ4, Ġ3, Ġ5, Ġ6, Ġ1—correspond to plausible numerical predictions that were actively suppressed. This supports the hypothesis that final layers may overwrite generalizable abstractions with memorized or noise signals.
These patterns are consistent with our mutual information analysis, which identifies intermediate layers as better semantically aligned with the prediction target—forming a ridge of generalization.
G.3 Implication
These decoding results reinforce our interpretation of the generalization ridge: intermediate layers contribute the most semantically informative updates to the model’s representation. The residual transitions thus serve as a useful lens for understanding how and where semantic meaning is introduced during forward propagation.
Appendix H Activation Patching Experiment
To assess which residual blocks causally support OOD prediction, we conduct a targeted activation-patching experiment. Let the model contain $L$ residual blocks with hidden states $Z_{\ell}\in\mathbb{R}^{d}$, and let $t^{*}$ denote the final context token (the last token before next-token prediction). For every OOD sequence $x_{\mathrm{ood}}$, we pair it with an ID sequence $x_{\mathrm{id}}$ sharing the same token structure. For each depth $\ell\in\{1,\dots,L\}$, we compute two forward passes: a standard OOD pass and a patched pass in which we replace only the hidden state of token $t^{*}$ at depth $\ell$ according to $Z_{\ell}(x_{\mathrm{ood}},t^{*})\leftarrow Z_{\ell}(x_{\mathrm{id}},t^{*}).$ All other activations remain unchanged. Let $logit_{\mathrm{base}}$ denote the logit assigned to the correct next token under the baseline OOD pass, and let $logit_{\mathrm{patch}}^{(\ell)}$ denote the corresponding logit under layer-$\ell$ patching. We measure causal impact using the logit drop $\Delta_{\ell}=logit_{\mathrm{base}}-logit_{\mathrm{patch}}^{(\ell)},$ where a positive value indicates that depth $\ell$ provides information necessary for correct OOD prediction.
Table 5 reports the averaged logit drop for all $\ell=1,\dots,12$ (GPT-2 Small on Synthetic Dataset). The effects are negligible for layers 1–8, begin rising around layer 9, increase sharply at layer 10, peak at layer 11, and decline again at layer 12. This rise–peak–fall pattern closely matches the ridge structure observed in Figure 2, where $I(Z_{\ell};Y)$ increases through mid-to-late depth, reaches a maximum in the ridge region, and then decreases. The activation-patching experiment therefore provides supporting causal evidence that the ridge layers identified by IndoRidge are necessary for successful OOD generalization.
Table 5: Average logit drop $\Delta_{\ell}$ (baseline logit - patched logit). Larger values indicate greater causal contribution to OOD prediction.
Layer $\ell$ 1 2 3 4 5 6 7 8 9 10 11 12
$\Delta_{\ell}$ -0.1037 -0.1039 -0.0327 -0.5405 -5.1663 -8.3992 -5.8914 -15.0996 -3.0998 9.3148 17.2022 15.5943
Appendix I Extending Predictive Information to Multi-Token Outputs
To complement our single-token analysis, we further evaluate information flow under multi-token outputs using the CNN/Daily Mail summarization dataset (1.0.0, test) (See et al., 2017; Hermann et al., 2015).
Autoregressive next-token information.
We follow the model’s natural decoding process. We analyze $I(Z_{\ell}^{(t)};\hat{Y}^{(t+1)})$ since there is no deterministic gold answer for each intermediate step in NLG, where for each generation step $t$, $Z_{\ell}^{(t)}$ is the hidden state of current token at layer $\ell$ and $\hat{Y}^{(t+1)}$ is the embedding of model’s predicted token at step $t+1$.
Across steps, the pattern is consistent (Figure 11): information rises from early layers, peaks in the middle layers, and then declines toward top layers. This gives a clear information ridge, indicating that intermediate layers encode the most stable and task-relevant core information that supports the model’s decision-making (generalization). In particular, high $I(Z_{\ell}^{(t)};\hat{Y}^{(t+1)})$ at these layers means the model’s predictive direction is most strongly determined by that layer: layers whose representations consistently constrain the next-token prediction across many contexts must encode the shared, invariant structure that the model relies on to generalize. In contrast, upper layers exhibit behavior that might be related to pretrained surface patterns from corpus such as token co-occurrence, causing the representation to drift away from this decision-relevant structure (memorization).
The first step is a notable outlier: it is the only step that conditions solely on the input prompt, and shows the sharpest and most pronounced ridge. Generating the first token requires a global transition from the prompt into the answer trajectory, which heavily engages the abstract, decision-relevant subspace in the middle layers. Crucially, this is also the point where information compression is most extreme, the model must condense all prompt-level semantics into a single initial token decision, therefore forms a substantial ridge. Once generation begins, later generation steps move into a more regular autoregressive regime, where each hidden state already contains a compressed summary of all previously generated content, and the continuation depends more on local context than on global restructuring. Consequently, the depth profiles become flatter and more stable across steps, although the mid-layer ridge remains visible.
(a) Qwen-2.5-0.5B on CNN/Daily Mail for 50 generation steps.
(b) LLaMA-3.1-8B on CNN/Daily Mail for 50 generation steps.
Figure 11: $I(Z_{\ell}^{(t)};\hat{Y}^{(t+1)})$ across layers (x-axis) and generation steps (y-axis), where the left y-axis shows first step mutual information (bars) and the right y-axis shows mutual information for the rest steps.
Pooled-answer information.
We quantify how much information each layer carries about the gold answer by computing $I(Z_{\ell}^{(t)};Y^{\text{pool}})$, where $Y^{\text{pool}}$ is a fixed target vector obtained by average-pooling the embedding-layer representations of all ground-truth answer tokens. Across models and steps (Figure 12), a consistent pattern emerges: intermediate layers are where the hidden state is most strongly aligned with the gold-answer.
(a) Qwen-2.5-0.5B on CNN/Daily Mail for 50 generation steps.
(b) LLaMA-3.1-8B on CNN/Daily Mail for 50 generation steps.
Figure 12: $I(Z_{\ell}^{(t)};Y^{\text{pool}})$ across layers (x-axis) and generation steps (y-axis), where the left y-axis shows first step mutual information (bars) and the right y-axis shows mutual information for the rest steps.
Taken together, the two constructions reveal a coherent picture of depth specialization in multi-token generation. Intermediate layers are simultaneously the most informative about the model’s next-token decisions and the most aligned with the global answer. Thus, the two observed ridges are complementary projections of a single underlying phenomenon: intermediate layers concentrated most meaningful and decision-relevant information. In contrast, upper layers increasingly reflect pretrained corpus-specific surface patterns, causing the representation to drift away from this decision-relevant structure.
(a) $I(Z_{\ell}^{(t)};\hat{Y}^{(t+1)})$: Qwen-2.5-0.5B on CNN/Daily Mail for 50 generation steps.
(b) $I(Z_{\ell}^{(t)};\hat{Y}^{(t+1)})$: LLaMA-3.1-8B on CNN/Daily Mail for 50 generation steps.
Figure 13: Mutual information comparing stopwords/punctuation against content words, showing that both stopwords/punctuation have consistently higher MI.
We also plot the mutual information comparing stopwords/punctuation against content words, showing that both stopwords/punctuation have consistently higher MI (Figure 13). These tokens are more predictable because they follow strict grammatical and structural patterns, whereas content words depend more on semantic context and therefore exhibit greater variability.
Appendix J Non-finetuned scenario
Beyond fine-tuning scenarios, we also computed predictive-information $I(Z_{\ell};Y)$ curves in a non-training zero-shot regime and observed the same mid-layer information peak pattern, see Figure 14.
Figure 14: GSM8K on Qwen2.5-0.5B shows an intermediate-layer MI rise and late-layer decline.
Appendix K Kernel Ablation
The matrix-based MI estimator requires a positive-definite Gram matrix, for which the Gaussian kernel is the standard choice. To assess robustness with respect to kernel selection, we additionally test the Laplacian and Polynomial kernels. While these kernels introduce quantitative shifts in the magnitude, the trend—including the rise–peak–decline structure and the location of the ridge—remains unchanged (Figure 15).
Polynomial Kernel.
Given a set of normalized representations $U=\{u_{i}\}_{i=1}^{N}$, the polynomial kernel Gram matrix is defined as
$(G^{\mathrm{poly}}_{U})_{ij}=\left(u_{i}^{\top}u_{j}+c_{0}\right)^{p},$ (4)
where $p$ is the polynomial degree and $c_{0}$ is a constant bias term.
Laplacian Kernel.
Given the same representation set $U=\{u_{i}\}_{i=1}^{N}$, the Laplacian kernel Gram matrix is defined as
$(G^{\mathrm{lap}}_{U})_{ij}=\exp\!\left(-\gamma\,\|u_{i}-u_{j}\|_{1}\right),$ (5)
where $\gamma>0$ is a kernel coefficient and $\|\cdot\|_{1}$ denotes the $\ell_{1}$ distance.
(a) Gaussian Kernel: Qwen-2.5-0.5B on ECQA.
(b) Polynomial Kernel: Qwen-2.5-0.5B on ECQA.
(c) Laplacian Kernel: Qwen-2.5-0.5B on ECQA.
Figure 15: Different kernel shows similar pattern.
Appendix L Use of Large Language Models.
Large language models (LLMs) were used solely as assistive tools for proofreading and improving clarity of writing.

### v3

@@ 7 Conclusion
We introduce InfoRidge, an information-theoretic framework designed to trace and quantify how information evolves across layers in transformer-based language models for natural language generation. By estimating both predictive information and incremental information gain, we systematically characterize the layerwise dynamics of information flow, offering a principled view of how signals are refined, amplified, or diminished as they propagate through the network. Our findings reveal a consistent generalization ridge emerging in intermediate layers, where mutual information between the hidden representation and the target label reaches its peak before gradually declining. This phenomenon reflects a fundamental trade-off between generalization and memorization as information flows deeper into the model. A similar pattern is observed in the multi-token setting. Attention analysis, capacity analysis and residual scaling experiments offer complementary causal evidence supporting this observation. Taken together, these findings position InfoRidge as a comprehensive framework for diagnosing how language models internally manage information during natural language generation, while also revealing the structural mechanisms that govern the balance between generalization and memorization.
Limitation
Our analysis considers a limited set of datasets and models. In addition, the kernel design for multi-token scenarios may not optimally capture the ground-truth distribution. Future work will extend analysis to broader model families and tasks and explore improved kernel formulations.
@@ A.1 Matrix-Based Entropy Estimation
Polynomial Kernel. Given a set of normalized representations $U=\{u_{i}\}_{i=1}^{N}$, the polynomial kernel Gram matrix is defined as $(G^{\mathrm{poly}}_{U})_{ij}=\left(u_{i}^{\top}u_{j}+c_{0}\right)^{p},$ where $p$ is the polynomial degree and $c_{0}$ is a constant bias term. Laplacian Kernel. Given the same representation set $U=\{u_{i}\}_{i=1}^{N}$, the Laplacian kernel Gram matrix is defined as $(G^{\mathrm{lap}}_{U})_{ij}=\exp\!\left(-\gamma\,\|u_{i}-u_{j}\|_{1}\right),$ where $\gamma>0$ is a kernel coefficient and $\|\cdot\|_{1}$ denotes the $\ell_{1}$ distance.
@@ Appendix B Dataset Overview and Statistics
Data Statistics
We conduct experiments on four datasets with varying levels of complexity and structure: CLUTRR (Sinha et al., 2019), ECQA (Aggarwal et al., 2021), CNN/DailyMail (See et al., 2017; Hermann et al., 2015) and a custom-designed Synthetic Arithmetic dataset. Table 3 summarizes key dataset statistics.
@@ B.2 ECQA
B.3 CNN/DailyMail
CNN/DailyMail (See et al., 2017; Hermann et al., 2015) is an summarization dataset containing news articles as written by journalists at CNN and the Daily Mail.
@@ C.1 Prompt Construction
All tasks are cast into a natural language generation format. The model receives a prompt and generate the next token. Below are construction strategies and examples for each dataset:
@@ E. Drive
CNN / DailyMail
[…]
Prompt:
(CNN)The Palestinian Authority officially became the 123rd member of the International Criminal Court on …
The following is a one-paragraph highlight summary of the article:
Target: Membership gives the ICC jurisdiction over alleged crimes committed in Palestinian territories since last June . Israel and the United States opposed the move, which could open the door to war crimes investigations against Israelis .
@@ C.2 Finetuning Settings
C.3 Residual Scaling with Learnable $\beta_{\ell}$ Parameters
@@
The in-distribution (ID) setting is using the training split of “gen_train23_test2to10”. Out-of-distribution (OOD) is conducted on held-out CLUTRR configurations that do not overlap with the training distribution: for GPT models we evaluate on the test split of “gen_train23_test2to10” (lengths 2 through 10), while for Qwen and LLaMA we use the test split of “rob_train_sup_23_test_all_23”, which defines a more challenging OOD setting by jointly varying both the clause length and the status of noise facts introduced into the story, given the higher modeling capacity of these models. In all cases, tasks 1.2 and 1.3 are excluded from the OOD sets, as they overlap with the training configuration and are therefore treated as in-distribution rather than out-of-distribution.
@@ Appendix D Compute and Licensing Details
CLUTRR (Sinha et al., 2019) is released under a CC BY 4.0 license as part of EMNLP 2019. The ECQA (Aggarwal et al., 2021) dataset is also released under a CC BY 4.0 license by its authors as part of EMNLP 2021. CNN/DailyMail is under Apache-2.0 License. The Synthetic Arithmetic dataset is custom-designed by the authors and does not rely on any external or licensed data sources.
@@ Appendix E Full Quantitative Results with Confidence Estimates
Predictive Information
We compute mutual information following the methodology in Appendix A. We subsample between 50 and 100 test examples per experiment and report results with statistical significance based on 5 seeds.
@@
Figure 9: Predictive information $I(Z;Y)$ across models and datasets. Lighter line colors represent later training epochs. Each curve shows the mean across 5 random seeds (0, 1, 2, 3, 42), and the shaded region denotes a 2-sigma (~96%) confidence interval.
In cases where the task is too simple relative to model capacity—such as the synthetic arithmetic task with LLaMA—this trend reflects an overfitting regime.
Incremental Information Gain
@@
Figure 10: Incremental information gain $I(\Delta Z;Y)$ across different models and datasets with ~96% CI error bars.
Kernel Ablation
To assess robustness with respect to kernel selection, we additionally test the Laplacian and Polynomial kernels (Figure 11).
(a) Gaussian Kernel: Qwen-2.5-0.5B on ECQA.
(b) Polynomial Kernel: Qwen-2.5-0.5B on ECQA.
(c) Laplacian Kernel: Qwen-2.5-0.5B on ECQA.
Figure 11: Different kernel shows similar pattern.
Residual Scaling
We present the complete set of residual scaling results, detailing the learned $\beta_{\ell}$ values across all transformer layers.
@@
Figure 12: Residual scaling coefficients $\beta_{\ell}$ across all transformer layers. Each curve shows the mean across five random seeds, and the shaded region denotes 1-sigma error bar. ID training emphasizes later layers, while OOD training has lower final layer weight. When the task is simple relative to the model’s capability, such as LLaMA on the synthetic task, the final-layer residual coefficients are similar for ID and OOD, consistent with Figure 9.